# General Riemann Integration on Partially Ordered Net Convergence Structures

### v1

@@ 6.1 Abstract Lebesgue Integration
    Define, then, using the last point, the integral of a function $f$ as the limit of the integrals of the approximating functions on $L$.
Nevertheless, a kind of general Lebesgue integral can be propose in this section if we let the uniqueness portion of the procedure be a part of the definition of this integral - something that can be checked in every case of interest (e.g, the references above).
This general integral is based on an extension of the sequential definition given in Section 3.2 of [12], and also in the ideas of Boccuto and Sambucini in [133] and [39].
In this subsection, we consider the following setting:
  1. 1.
    $(X,\eta_{1}),(Y,\eta_{2}),(Z,\eta_{3})$ three spaces equipped with Boccuto-Candeloro net convergence structures, and connected by a product as in Definition 3.6,
  2. 2.
    $(\Omega,\mathcal{H})$ an abstract paved space,
  3. 3.
@@
    $\mu:\mathcal{H}\rightarrow Y^{+}$ a finitely additive set function.
Now, to begin the procedures to define the integral, consider a space of functions $L$, consisting of all $f:\Omega\rightarrow X$ satisfying some fixed property (e.g, being simple, bounded, etc) such that it contains all simple functions defined on $(\Omega,\mathcal{H})$, and with the property that $f\mathbbm{1}_{A}\in L$, for each $A\in\mathcal{H}$.
Suppose now that there exists a functional denoted by:
$\int_{\Omega}^{L}(\cdot)\ d\mu:L\rightarrow Z,$
such that:
  1. 1.
    $\int^{L}_{\Omega}(\cdot)\ d\mu$ is linear,
  2. 2.
    $\int^{L}_{\Omega}(\cdot)\ d\mu$ is such that, for each simple function $g=\sum_{i=1}^{n}a_{i}\mathbbm{1}_{A}$,
    $\int^{L}_{\Omega}gd\mu=\sum_{i=1}^{n}a_{i}\mu(A_{i}),$
Now put $\int^{L}_{A}fd\mu=\int^{L}_{\Omega}f\mathbbm{1}_{A}d\mu$.
Definition 6.1 (Abstract Lebesgue Integral).
A non-negative function $f:\Omega\rightarrow X$ is said to be Abstractly Lebesgue integrable, or just Lebesgue integrable in case no confusion arises, if there is a net $\{f_{i}\}_{i\in I}$ of functions in $L$ such that $f_{i}\rightarrow f$ in some convergence defined on the function space $L$ with respect to $(X,\eta_{1})$ [^7] For example: the uniform or almost uniform convergences defined in the previous sections., which we call the approximation procedure, and such that the limit:
$\lim_{(I,\eta_{3})}\int^{L}_{A}f_{i}d\mu,$
$I_{f}(A)=\lim_{(I,\eta_{3})}\int^{L}_{A}f_{i}d\mu,$
the Abstract Lebesgue integral of $f$ on $A$.
In this case, an arbitrary function $f:\Omega\rightarrow X$ is called Abstractly Lebesgue integrable if $f^{+}$ and $f^{-}$ are and in this case the integral of $f$ is defined to be:
$I_{f}(A)=I_{f^{+}}(A)-I_{f^{-}}(A),$
where $A\in\mathcal{H}$.
The first thing to notice is that this integral is readily seen to be a linear operator on the space of Lebesgue integrable functions. Other properties of this integral seem to require special hypothesis, such as the uniform integrability property of the net approximation of functions as in [12] or some other restrictive property of the approximation procedure.
We also comment that not only the sequential (general) Lebesgue integral of [12] is contained in our definition, but also most examples of the Lebesgue type integrals cited in the beginning of Section 5. In particular, taking $L$ to be the space of elementary functions, and approximation procedure of almost uniform convergence, we get the Pavlakos Integral. Also, by analogous constructions, the recently defined integrals of [41], [42], defined on the basis of a sequential (monotone) approximation by simple functions, are also contained in our general definition. Lastly, by the results in Section 5, we get some some connections between this general Lebesgue Integral and the Riemmanian ones.
We stop the development of this integral at this point, and refer the reader to [12], [133] and [39] for more properties in some specific cases, and also point that some other questions will be answered in a future work.
6.2 A Saks Type Integral
@@
Now, we pass to a brief discussion of the last type of integral to be defined in this work, a kind of Saks type integral as defined originally by Saks [134] and Gould [135], but also more recently by Fleischer [15] and Shcherbin [19]. Basically, the idea is to integrate functions by first integrating in an upwards directed class of sets, and then take the limit of integration in sets in this class as it "expands" (as it is (upwards) directed by inclusion, therefore forming a net). The idea is the same as in classical Lebesgue integration, where we may define the integral in an unbounded set as the limit of integrals on approximating (upwards) bounded sets.
For this section, we mainly follow the ideas of Fleischer [15] (which are, interestingly, presented in a convergence structure type language), and the same structures as in the Definition 3.9 of the Fleischer Net Riemann integral.
In this case, we have, by the same notation as in Section 3.1 of the (Fleischer) Net Riemann Integral:
@@
and $f$ is Fleischer Net Riemann integrable on each $B\in\mathcal{C}_{f,A}$. Then, we say that $f$ is Saks integrable on $A$ if the net of Fleischer Net Riemann integrals $\{\int_{B}fd\mu\}_{B\in\mathcal{C}_{f,A}}$ has a limit in $Z$ and, in this case, we define:

### v2

@@ Abstract
    $\mu$ is of finite $\sigma_{\phi}$-finite $(\mathcal{U},\mathcal{W})$-semivariation if there exists a $\phi\in\Phi$ such that for every $U\in\mathcal{U}$ the operator valued measure $\mu$ is of $\sigma_{\phi}$-finite $(\mathcal{U},\mathcal{W})$-semivariation.
  4. (iv)
    $\nu:\sigma(\mathcal{H})\rightarrow Y$ finite additive is $(W,\sigma)$-additive, for $W\in\mathcal{W}$, if $\nu$ is countable additive in $Y_{W}$. We then say that $\nu$ is $(\mathcal{W},\sigma)$-additive if there exists $W\in\mathcal{W}$ such that $\nu$ is $(W,\sigma)$-additive.
  5. (v)
    Let $\phi\in\Phi$. We say that $\mu$ is $\sigma_{\phi}$-additive if $m$ is of finite $\sigma_{\phi}$-finite $(\mathcal{U},\mathcal{W})$-semivariation and, for every $A\in\mathcal{H}_{U,\phi(U)}$, the finite additive set function $\mu(A\cap\cdot)x:\sigma(\mathcal{H})\rightarrow Y$ is a $(\phi(U),\sigma)$-additive for every $x\in X_{U}$, $U\in\mathcal{U}$. We then say that $\mu$ is $\sigma_{\Phi}$-additive if there exists $\phi\in\Phi$ such that $\mu$ is $\sigma_{\phi}$-additive.
  6. (vi)
    Let $W\in\mathcal{W}$. A sequence of $(W,\sigma)$-additive set functions $\nu_{n}:\sigma(\mathcal{H})\rightarrow Y$, $n\in\mathbb{N}$, is said to be uniformly $(W,\sigma)$-additive on $\sigma(\mathcal{H})$ if for every $\varepsilon>0$, $A\in\sigma(\mathcal{H})$ such that $p_{W}(\nu_{n}(A))<\infty$, $A_{i}\in\sigma(\mathcal{H})$ disjoint, there exists $k_{0}\in\mathbb{N}$ such that for every $k\geq k_{0}$,
    $p_{W}(\nu_{n}(\bigcup_{i=k+1}^{\infty}(A_{i}\cap E)))\leq\varepsilon,$
    uniformly for every $n\in\mathbb{N}$.
The last step to define the integral is to define some structures related to a class of "measurable" functions in this abstract (operator valued) setting. We again collect the relevant definitions from Duchoň and Haluška (2005):
Definition 5.8 (Measurable Functions and Bornologies).
@@
For $A\in\mathcal{H}$ and $\mu:\mathcal{H}\rightarrow L(X,Y)$ a finitely additive operator valued measure, we say that:
  1. (i)
    For $R\in\mathcal{U}$, a sequence of functions $f_{n}:\Omega\rightarrow X$, $n\in\mathbb{N}$, $(R,A)$-converges $\hat{\mu}_{U,W}$-a.e. to a function $f:\Omega\rightarrow X$ if:
    $\lim_{n\rightarrow\infty}p_{R}(f_{n}(\omega)-f(\omega))=0,$
    for every $\omega\in A\setminus N$, where $N$ is a $\hat{\mu}_{U,W}$-null set.
  2. (ii)
    A sequence of functions $f_{n}:\Omega\rightarrow X$, $n\in\mathbb{N}$, $(\mathcal{U},A)$-converges $\hat{\mu}_{\mathcal{U},\mathcal{W}}$-a.e. to a function $f:\Omega\rightarrow X$ if there exists $R\in\mathcal{U}$ and $(U,W)\in\mathcal{U}\times\mathcal{W}$ such that the sequence $f_{n}$ $(R,A)$-converges $\hat{\mu}_{U,W}$-a.e. to f. In this case, we write $f=\mathcal{U}-\lim_{n\rightarrow\infty}f_{n}$ $\hat{\mu}_{U,W}$-a.e.
  3. (iii)
    We say that a function $f:\Omega\rightarrow X$ is $\mathcal{U}$-measurable function if there exists a $R\in\mathcal{U}$ such that for every $R\subseteq U\in\mathcal{U}$ and $\delta>0$
    $\{\omega\in\Omega:p_{U}(f(\omega))\geq\delta\}\in\sigma(\mathcal{H}).$
    We denote by $\mathbb{M}_{\mathcal{U}}$ the set of all $\mathcal{U}$-measurable functions.
  4. (iv)
    A function $f:\Omega\rightarrow X$ is $\mathcal{H}$-simple if it has a finite range and $f^{-1}(x)\in\mathcal{H}$ for each $x\in X\setminus\{0\}$. We denote the space of all $\mathcal{H}$-simple functions by $\mathcal{S}$. On the other hand, we say that a function $f:\Omega\rightarrow X$ is $\mathcal{H}_{U,W}$ simple, for $(U,W)\in\mathcal{U}\times\mathcal{W}$ if:
    $f=\sum_{i=1}^{n}x_{i}\mathbbm{1}_{A_{i}},$
    where $\{x_{i}\}_{i=1}^{n}\in X_{U}$, $\{A_{i}\}_{i=1}^{n}$ is a disjoint sequence of sets in $\mathcal{H}_{U,W}$ and $n\in\mathbb{N}$. We denote the space of these functions by $\mathcal{S}_{U,W}$. Then, a function is said to be $\mathcal{H}_{\mathcal{U},\mathcal{W}}$ simple if there exists $(U,W)\in\mathcal{U}\times\mathcal{W}$ such that $f\in\mathcal{S}_{U,W}$. The space of all $\mathcal{H}_{\mathcal{U},\mathcal{W}}$ simple functions will be denoted by $\mathcal{S}_{\mathcal{U},\mathcal{W}}$.
We now give the main definition of the integral first presented in Haluška (1988), which is a generalization of the Banach space operator-valued measure of Dobrakov (1970):
Definition 5.9 (Haluska Integral).
Let $\mu:\mathcal{H}\rightarrow L(X,Y)$ be a $\sigma_{\phi}$-additive operator valued measure. A function $f\in\mathbb{M}_{\mathcal{U}}$ is said to be $\mathcal{H}_{\mathcal{U},\mathcal{W}}$-integrable, or Haluska integrable, and we write $f\in\mathcal{I}_{\mathcal{U},\mathcal{W}}$, if:
  1. (i)
    There exists a sequence $f_{n}\in\mathcal{S}_{\mathcal{U},\mathcal{W}}$, $n\in\mathbb{N}$, of (simple) functions such that:
    $\mathcal{U}-\lim_{n\rightarrow\infty}f_{n}=f\ \hat{\mu}_{\mathcal{U},\mathcal{W}}-a.e.,$
  2. (ii)
    The integrals (defined, as usually, additively) $\int_{A}f_{n}d\mu$, $n\in\mathbb{N}$, are uniformly $(\mathcal{W},\sigma)$-additive measures on $\sigma(\mathcal{H})$.
Then, the (Haluska) integral of the function $f\in\mathcal{I}_{\mathcal{U},\mathcal{W}}$ on a set $A\in\sigma(\mathcal{H})$ is defined by the equality:
$\int_{A}^{(H)}fd\mu=\mathcal{W}-\lim_{n\rightarrow\infty}\int_{A}f_{n}d\mu.$
The $S^{*}$-partition integral of Definition 3.6 takes the following form in this context:
Definition 5.10 ($S^{*}$-Partition Integral and Bornologies).
Let $(U,W)\in\mathcal{U}\times\mathcal{W}$. A function $f:\Omega\rightarrow X$ is $S^{*}_{U,W}$-partition integrable on $A\in\mathcal{H}$ if, for $f_{U}:\Omega\rightarrow X_{U}$, there exists $y\in Y_{W}$ such that for every $\varepsilon>0$ there exists a $P_{0}\in\mathcal{P}(A)$ consisting of $\mathcal{H}_{U,W}$ sets such that, for every $P\in\mathcal{P}(A)$ consisting of $\mathcal{H}_{U,W}$ sets and which satisfies $P\geq P_{0}$, the series $\sum_{\alpha\in P}\mu(\alpha)f_{U}(\delta_{\alpha})$ is unconditionally convergent in $Y_{W}$ for every choice function $\delta$ and:
$p_{W}(\sum_{\alpha\in P}\mu(\alpha)f_{U}(\delta_{\alpha})-y)\leq\varepsilon.$
We then have the main result of the article Duchoň and Haluška (2005), which gives the equivalence between the $S^{*}$-partition integral above and the Haluska integral, and which generalizes the corresponding result of Dobrakov (1988) for the Banach-valued case:
Theorem 5.5 ($S^{*}$-Partition Integral is Equivalent to the Haluska Integral).
A $\mathcal{H}_{\mathcal{U},\mathcal{W}}$ measurable function $f:\Omega\rightarrow X$ is Haluska integrable on $\sigma(\mathcal{H})$ if and only if it is $S^{*}$-partition integrable on $\sigma(\mathcal{H})$ and, in this case,
$\int_{A}^{(H)}fd\mu=\int_{A}^{S^{*}}fd\mu.$
Using the class of Dobrakov measurable functions, which is a special class of the class of measurable functions defined above, we get that, from the last theorem, $X,Y$ being Banach spaces,
Corollary 5.5.1 ($S^{*}$-Partition Integral is Equivalent to the Dobrakov Integral).
A Dobrakov measurable function $f:\Omega\rightarrow X$ is Dobrakov integrable on $\sigma(\mathcal{H})$ if and only if it is $S^{*}$-partition integrable on $\sigma(\mathcal{H})$ and, in this case,
$\int_{A}^{(D)}fd\mu=\int_{A}^{S^{*}}fd\mu,$
where $\int_{A}^{(D)}fd\mu$ is de Dobrakov integral of $f$ on $A$.
These two equivalences results gives interesting further properties of the net Riemann integral, and its special (topological) case of the $S^{*}$-partition integral, in the context of bilinear integration in locally convex spaces, specially Banach and bornological ones. For further interesting convergence theorems results, we refer the reader to the articles cited above.
We should also comment that, as noticed by Haluška (1997), his integral contains the Egorov type integral of the Spanish school constructed by Rodríguez Salazar (1985), thus giving another interesting integration procedure related to the net Riemann integral. Interestingly, we also may have proved this result directly by looking at the constructions in Rodríguez Salazar (1985) or Fernández y Fernández-Arroyo et al. (1995) and noticing that it reduces to approximating functions by countable valued functions in the regime of quasi-uniform convergence (always supposing finite semivariation of the operator valued-measures involved), thus the equivalence can be proven using Theorem 4.12. We don’t pursue this construction for brevity.
5.2.3 The Submeasure Integral of Massé.
We now study the last point in the integration equivalences. This is the integral of Masse (1979), which contains as special case the integral of Marik (1987). We mainly follow the first cited article, and we also use the submeasure terminology originally due to Drewnowski (1972a).
For this section, unless stated to the contrary, we consider $X,Y,Z$ Hausdorff topological vector spaces each of which has the net convergence structure induced by its topology, which we assume are complete and generated by a base of balanced zero-neighborhoods, denoted in each case by $\mathcal{U}_{X},\mathcal{U}_{Y},\mathcal{U}_{Z}$. Also, $Fs(X)$ well denote any base of continuous $F$-seminorms which define the topology of $X$. In this case, with a bilinear product, we fix an integration structure given by Definition 2.3.
Also, in this section, $\mathcal{H}$ will be an algebra of subsets of $\Omega$, and we denote by $\Delta$ the operation of symmetric differences on sets. In this case, we define:
$\mathbb{S}=\{\xi_{i}\}_{i\in I},$
to be a non-empty collection of (extended) real valued submeasures defined on $\mathcal{H}$. In this case, we equip $(2^{\Omega},\Delta,\cap)$ with the topology induced by $\mathbb{S}$ for which it is a topological ring (with operations of symmetric difference and intersection). We call this topology $\tau_{\mathbb{S}}$. Then, $\mathcal{H}(\mathbb{S})$ will denote the sub-ring of this last topological ring with the subspace topology.
We now define an essential concept due to Masse (1979), which generalizes the idea of convergence in measure:
Definition 5.11 (Convergence in a Family of Submeasures).
We say that a net of functions $\{f_{l}\}_{l\in L}$, each defined on $\Omega$ and with values in $X$, converges in submeasure to $f:\Omega\rightarrow X$ if and only if:
[…]
for each $d\in Fs(X)$ and $i\in I$, where $||\cdot||_{d}$ is the $F$-norm induced on $X$ by $d$. The (vector [^23] See section 1 of Masse (1979).) topology induced by this convergence in $X^{\Omega}$, which has the base of neighborhoods of the form $\{W_{K}(U,\varepsilon)\}_{U\in\mathcal{U}_{X},K\subseteq\mathcal{F}(\mathbb{N}),\varepsilon>0}$, where:
$W_{K}(U,\varepsilon)=\{g\in X^{\Omega}:\xi_{k}(\{\omega\in\Omega:f-g\notin U\})\leq\varepsilon,\ k\in K\},$
is called the topology of convergence in $\mathbb{S}$-submeasure.
We now define a structure that consists, in a generalized sense, on the sets for which we may apply the family $\mathbb{S}$ (one example would be the sets of finite semivariation in the previous section):
Definition 5.12 ($\mathbb{S}_{\mu}$-Integrable Sets).
For a finitely additive set function $\mu:\mathcal{H}\rightarrow Y$ and the family of submeasures $\mathbb{S}$, we suppose the existence of a sub-ring of $\mathcal{H}$, which we denote by $\mathcal{H}^{\mu}_{\mathbb{S}}$, and call the family of $\mathbb{S}_{\mu}$-integrable sets.
Associated to these structures, we have the main two hypothesis of Masse (1979), which are imposed in order to give some regularity to the integration procedure, and we assume for the rest of the section. Basically, they ensure the to-be-defined integration procedure is continuous in the space of simple functions with uniform convergence:
Definition 5.13 (Massé’s Hypothesis).
We suppose that the integration structure of this section satisfies the following two hypothesis:
($M_{1}$) The integration operator acting on simple functions, based on $\mathcal{H}^{\mu}_{\mathbb{S}}$,
$\int_{\Omega}\cdot\ d\mu:(S^{\mathcal{H}^{\mu}_{\mathbb{S}}}_{\infty},\tau_{\infty})\rightarrow(Z,\tau),$
given by:
$\int_{\Omega}sd\mu=\sum_{i=1}^{n}s_{i}\mu(A_{i}),$
is continuous, where $(S^{\mathcal{H}^{\mu}_{\mathbb{S}}}_{\infty},\tau_{\infty})$ is the space of simple functions, based on $\mathcal{H}^{\mu}_{\mathbb{S}}$, with the topology of uniform convergence induced on $X^{\Omega}$ denoted by $\tau_{\infty}$.
($M_{2}$) For every $x\in\mathcal{X}$,
$\lim_{\mathcal{H}^{\mu}_{\mathbb{S}},E\rightarrow\emptyset}x\mu(E)=0,$
where the limit is taken in $\mathcal{H}^{\mu}_{\mathbb{S}}$ equipped with the convergence in $\mathbb{S}$-submeasure topology induced by indicator function convergence.
We now give the main integral of Masse (1979):
Definition 5.14 (Massé Integral).
A function $f:\Omega\rightarrow X$ is said to be Massé-integrable with respect to a finitely additive set function $\mu:\Omega\rightarrow Y$ if there exists a net of $X$-valued simple functions based on $\mathcal{H}^{\mu}_{\mathbb{S}}$, denoted by $\{f_{l}\}_{l\in L}$, such that $f_{l}\rightarrow f$ in $\mathbb{S}$-submeasure and such that the net of additively-defined integrals $\{\int_{A}^{(M)}f_{l}d\mu\}_{l\in L}$ is Cauchy in $Z$ uniformly for $E\in\mathcal{H}$ .
In this case, we define the Massé-integral of $f$ on $A\in\mathcal{H}$ by:
$\int_{A}^{(M)}fd\mu=\lim_{l\in L}\int_{A}^{(M)}f_{l}d\mu.$
We refer the reader to section 2 of Masse (1979) for the facts that this integral is well-defined, is an homomorphism in the space of integrable functions, and also for Vitali type convergence theorems.
[…]
Example 5.1 (Sion Integrable, but not Massé Integrable, Function).
Let $\Omega=\mathbb{N}$, $\mathcal{H}$ the algebra of cofinite sets in $\Omega$, and take $\mu(A)$ (the finitely, but not $\sigma$,additive set function) to be 0 when $A$ is finite, and $1$ when $A^{c}$ is finite. Then, it is straightforward to see that the function $f=\mathbbm{1}_{\mathbb{N}}$ is Sion integrable, with integral 0, and Massé integrable, with integral $1$.
In fact, this example gives an indication of a much wider phenomena, as can be seen from the next proposition, which is a consequence of Theorems 3.8 and 3.1:
Proposition 5.9.
Let $f$ be a Massé and $S^{*}$/Sion-integrable function with respect to a finitely additive measure $\mu:\mathcal{H}\rightarrow Y$ such that $\nu_{f}:\mathcal{H}\rightarrow Y$, the set function induced by the Massé integral of $\mathcal{H}$, is not $\sigma$-additive on $\mathcal{H}$. Then, the Massé and $S^{*}$/Sion-integrals of $f$ are not equal.
In other words, from the fact that the $S^{*}$/Sion-partition integrals always project a measure intro its $\sigma$-additive part we may, for example, picking a purely finitely-additive set function, always produce functions and set functions for which these integrals disagree.
Therefore, we need further hypothesis, as in the sections above, to prove some equivalence results concerning these integrals, or some other interesting versions of the net Riemann integral. Nevertheless, we couldn’t find a single general condition on the Sion or $S^{*}$-integrals, for example, that gives equivalence with Massé’s one, something that is difficult given the radically different "natures" of these two integrals, as argued by Massé in the epilogue of his thesis (see Massé (1974)), and in the penultimate paragraph of Section 2 of Masse (1979), as well as by the example and result above.
One possible approach for the resolution of this problem may be to reduce it to a question of which type of convergence in submeasure theorem for the Net Riemann integral, or its special cases, is valid. As said in our section about convergence theorems, this is left for a future work, and we therefore leave the general version of the equivalence as an open problem.
The second approach, without asking the $\sigma$-additivity of $\mu$, may be to look for integrals in which the projection property into the $\sigma$-additive part is not automatic, and restrict the class of functions for which some type of convergence theorem is valid. In fact, by $(M1)$ we know that every function which is a uniform limit of a net of simple functions based $\mathcal{H}^{\mu}_{\mathbb{S}}$ is Massé integrable. To follow this approach, we need the next result, which we state for uniform spaces (see Page (1988), or the other topology references above, for non-defined terms):
Lemma 5.2 (Uniform Approximation in Totally Bounded Uniform Spaces).
Let $f:\Omega\rightarrow X$ be a map from a paved algebra $(\Omega,\mathcal{H})$ to a totally bounded uniform space $(X,\mathcal{U})$. Then, $f$ is the uniform limit of a net $\{f_{i}\}_{i\in I}$ of simple functions $f_{i}:\Omega\rightarrow X$. If $f^{-1}(U[\cdot])\in\mathcal{H}$ for every $U\in\mathcal{U}$  [^24] Or a (sub)base for this uniformity, which is equivalent., then we can choose the net $\{f_{i}\}_{i\in I}$ to be measurable with respect to $\mathcal{H}$.
Proof.
As $X$ is totally bounded, for each $U\in\mathcal{U}$ there exists a finite collection of points $\{x_{k}\}$, $k=1,\dots,n_{U}$, $n_{U}\in\mathbb{N}$ - which depends on the entourage - such that:
$X=\bigcup_{k=1}^{n_{U}}U(x_{k}).$
Now, let $A_{k}^{U}=f^{-1}(U(x_{k})-\bigcup_{j=1}^{k-1}U(x_{j}))$, for $k=2,\dots,n_{U}$, and $A_{1}^{U}=f^{-1}(U(x_{1}))$. Then, for each (fixed) $U\in\mathcal{U}$, the sets $A_{k}^{U}$ are disjoint and $\Omega=\bigcup_{k=1}^{n_{U}}A_{k}^{U}$. Therefore, define the simple function given by:
$f_{U}(\omega)=\sum_{k=1}^{n_{U}}x_{k}\mathbbm{1}_{A_{k}^{U}}(\omega),$
which, by definition, have $f(\omega)\in U(f_{U}(\omega))$, for each $U\in\mathcal{U}$, and $\omega\in\Omega$. Therefore, by definition, we have the uniform convergence $f_{U}\rightarrow f$, when $\mathcal{U}$ with the subset (downwards!) directed order, and therefore $\{f_{U}\}_{U\in\mathcal{U}}$ being a well defined net. Also, when $f$ satisfies the preimage condition $f^{-1}(U)\in\mathcal{H}$ for every $U\in\mathcal{U}$, then the support sets of each $f_{U}$ are in $\mathcal{H}$, and therefore all of the approximation maps are measurable with respect to the algebra $\mathcal{H}$. ∎
Therefore, from this last result, and the fact that the $S$-Kolmogorov integral (see Example 2.1) agrees with the Massé integral on simple functions and satisfy the hypothesis of the (uniform convergence) Theorem 4.13, we have:
Corollary 5.5.2 (Massé and $S$-Kolmogorov Integrals).
Suppose that $f:\Omega\rightarrow X$ is a totally bounded function such that $f^{-1}(O)\in\mathcal{H}^{\mu}_{\mathbb{S}}$ for each open set $O\subseteq X$. Then, $f$ is $S$-Kolmogorov and Massé integrable, and the two integrals agree.
For a set of minimal, but strong, hypothesis, this seems the most general result we have without supposing further hypothesis on the additivity of $\mu$.
If we allow $\mu$ to be $\sigma$-additive, then is is readily seen that we can change the integral in Corollary 5.5.2 to be the $S^{*}$/Sion integral and get the same result. Also, in this case, we now study a different set of restrictive hypothesis that make the referenced integrals equal . For that, we first need (Massé, 1974, Lemma 6.11, p. 30), and then (Massé, 1974, Theorem 9.7, p. 60), combined with (Massé, 1974, Theorem 6.4-6.6, p. 27):
Theorem 5.6.
Let $\mathbb{S}$ be a non-empty family of (extended real) submeasures in a $\sigma$-algebra $\mathcal{H}$, such that it possesses a control submeasure [^25] For this definition, see Drewnowski (1974).. Suppose also that $X$ and $Z$ are metrizable. Then, a function $f:\Omega\rightarrow X$ is Massé integrable, with respect to the family $\mathbb{S}$, if and only if there is a sequence of simple functions based on $\mathcal{H}^{\mu}_{\mathbb{S}}$ such that:
  1. (i)
    There exists a countable collection of subsets of $\Omega$, all elements of $\mathcal{H}$, $\{F_{k}\}_{k\in\mathbb{N}}$, such that $\Omega\setminus\bigcup_{k\in\mathbb{N}}F_{k}$ is a $\mu$-null set and $f_{n}\rightarrow f$ uniformly on each $F_{k}$;
  2. (ii)
    $\{\int_{E}^{(M)}f_{n}d\mu\}_{n\in\mathbb{N}}$ is Cauchy in $Z$ uniformly for $E\in\mathcal{H}$.
In this case, we shall modify the proof of the main theorems of Duchoň and Haluška (2005) and Dobrakov (1988) to prove the following:
Theorem 5.7.
Assume the hypothesis of Theorem 5.6, and let $f:\Omega\rightarrow X$ be Massé integrable on each $A\in\mathcal{H}$ with respect to $\mu:\mathcal{H}\rightarrow Y$, which is $\sigma$-additive. Suppose also $Y$ is metrizable and the semivariation of $\mu$ is finite on each $\mathcal{H}$. Then, $f$ is $S^{*}$-integrable on each $A\in\mathcal{H}$, and the two integrals agree.
Proof.
We begin by noting that, by hypothesis, we may, using Theorem 5.6, find a countable collection of subsets of partition of $\Omega$, based on sets from $\mathcal{H}$, $\{F_{k}\}_{k\in\mathbb{N}}$, such that $\Omega\setminus\bigcup_{k\in\mathbb{N}}F_{k}$ is a $\mu$-null set and $f_{n}\rightarrow f$ uniformly on each $F_{k}$, which are each of finite semivariation and may be assumed, without loss of generality, to be increasing. Also, by the fact that the Massé integral ignores null sets in the sense of Observation 5.1 (see Masse (1979)), we may reduce the problem to the integrability and equality of integrals of $f\mathbbm{1}_{\Omega\setminus(\bigcup_{k\in\mathbb{N}}F_{k})^{c}}$, which we shall denote by $g$.
By the uniform convergence theorem given in Theorem 4.13, which is applicable as we assume $(M1)$, $f$ is $S^{*}$ and $S$-integrable on each $F_{k+1}\setminus F_{k}$, as they are uniform limits of simple functions on these sets and the set function $\mu$ is $\sigma$-additive (therefore, each simple function is $S^{*}$ integrable), and the integral equals the Massé integral by the same argument before Corollary 5.5.2. Now using these two facts we may, for $\varepsilon>0$ and $k\in\mathbb{N}$ both fixed, find a finite partition $P_{0}\in\mathcal{P}^{f}(F_{k+1}\setminus F_{k})$ such that for each countable partition $P_{0}^{(k)}\in\mathcal{P}(F_{k+1}\setminus F_{k})$ that refines $P_{0}$, and any choice function $\delta$, the series $\sum_{\alpha\in P_{0}^{(k)}}g(\delta_{\alpha})\mu(\alpha)$ is unconditionally summable and:
$|\sum_{\alpha\in P_{0}^{(k)}}g(\delta_{\alpha})\mu(\alpha)-\int^{(M)}_{A\cap F_{k+1}\setminus F_{k}}gd\mu|\leq\frac{\varepsilon}{2^{k+1}},$ (31)
for each $A\in\mathcal{H}$ and $|\cdot|$ a fixed element of $Fs(Z)$.
Now, put $P_{0}^{F}=\bigcup_{k\in\mathbb{N}}P_{0}^{(k)}$, which is an element of $\mathcal{P}(F)$. Take now $P\geq P_{0}^{F}$ in $\mathcal{P}(F)$, and notice that $P$ is of the form:
$P=\bigcup_{k\in\mathbb{N}}P^{(k)},$
for $P^{(k)}\in\mathcal{P}(F_{k+1}\setminus F_{k})$ and $P^{(k)}\geq P_{0}^{(k)}$. We now show that:
$\sum_{\alpha\in P}g(\delta_{\alpha})\mu(\alpha)=\sum_{k\in\mathbb{N},\alpha^{(k)}\in P^{(k)}}g(\delta_{\alpha^{(k)}})\mu(\alpha^{(k)}),$
is unconditionally summable. In fact, it is sufficient to stablish that, by Antosik interchange theorem (see (Swartz, 1996, Theorem 1, Section 8.5, p. 98)):
  1. (i)
    For each $k\in\mathbb{N}$, the series $\sum_{\alpha\in P^{(k)}}g(\delta_{\alpha})\mu(\alpha)$ is unconditionally summable in $Z$. This holds because $g$ is $S^{*}$ integrable on each $F_{k+1}\setminus F_{k}$ by the argument above.
  2. (ii)
    For any sequence of subsets $I^{(k)}\subseteq\mathbb{N}$, $k\in\mathbb{N}$, the series:
    $\sum_{k=1}^{\infty}\sum_{j\in I^{(k)}}g(\alpha^{(k)}_{j})\mu(\alpha^{(k)}_{j})$,
    is unconditionally summable in $Z$.
For the second point, note that, for each fixed $I^{(k)}$ (as above) and by 31,
$|\sum_{j\in I^{(k)}}g(\alpha^{(k)}_{j})\mu(\alpha^{(k)}_{j})-\int^{(M)}_{\bigcup_{j\in I^{(k)}}\alpha^{(k)}_{j}\cap F_{k+1}\setminus F_{k}}gd\mu|\leq\frac{\varepsilon}{2^{k+1}},$
for each $k\in\mathbb{N}$ and $|\cdot|\in Fs(Z)$ fixed as before. Also, by using this inequality and the countable additivity of the Massé integral on $\mathcal{H}$ (as $\mu$ is countably additive - see Proposition $2.6$ of Masse (1979)), we get that
$\sum_{k\in\mathbb{N}}\int^{(M)}_{\bigcup_{j\in I^{(k)}}\alpha^{(k)}_{j}\cap F_{k+1}\setminus F_{k}}gd\mu,$
is unconditionally summable in $Z$. By this result and using again inequality 31, we get that:
$|\sum_{\alpha\in P}g(\delta_{\alpha})\mu(\alpha)-\int_{A}gd\mu|\leq\varepsilon,$
with the same notation as before. Therefore, by definition, as $|\cdot|\in Fs(Z)$ was arbitrary, $g$ is $S^{*}$ integrable on $A\in\mathcal{H}$ and this integral equals the Massé integral of $f$ on $A$. ∎
With this result, we finish the main points of equivalence regarding Lebesgue type integration and the net Riemann integral in topological vector spaces. We now pass to a discussion of another (the last) type of integration.
6 A Saks Type Integration Procedure in Net Convergence Structures.
@@
Now, we pass to, as indicated in the previous section, a brief discussion of the last type of integral to be defined in this work, a kind of Saks type integral as defined originally by Saks (1964) and Gould (1965), but also more recently by Fleischer (1990) and Shcherbin (1998), all in topological or partially ordered cases. Basically, the idea is to integrate functions by first integrating in an upwards directed class of sets, and then take the limit of integration in sets in this class as it "expands" (as it is (upwards) directed by inclusion, therefore forming a net). This concept is a classical procedure in classical Lebesgue integration, where we may define the integral in an unbounded set as the limit of integrals on approximating (upwards) bounded sets.
An interesting aspect of this integration procedure is that, by a Riemmanian definition, we may integrate with respect to non-additive set functions and get, by the set function induced by this process, a non-additive set function. As seen in Section 2, this is not the case for the net Riemann integral, which causes problems when studying the relation between it and integration procedures of Lebesgue type against non-additive set functions, such as the ones contained in the classical reference by Pap (1995).
Considering the same integration structures and notation as in Section 2, we may directly define the main integral of this section:
@@
and $f$ is net Riemann integrable on each $B\in\mathcal{C}_{f,A}$. Then, we say that $f$ is Saks integrable on $A$ if the net of net Riemann integrals $\{\int_{B}fd\mu\}_{B\in\mathcal{C}_{f,A}}$ has a limit in $Z$ and, in this case, we define: